OpenStudy (omnomnom):

How do I do this? I sorta forgot so yeah... x + y = 3x - 2y

OpenStudy (omnomnom):

I need to find both x and y

OpenStudy (anonymous):

by grouping (xy-3x)(-2y+6) x(y-3)-2(y-3) (x-2)(y-3)^2

OpenStudy (omnomnom):

What. O.o

OpenStudy (anonymous):

does it helpsu or not

OpenStudy (omnomnom):

Um... Not really i dont think i learned that.. But idk i forgot

OpenStudy (anonymous):

best response

OpenStudy (aripotta):

no idea :o are you trying to find x in terms of y?

OpenStudy (aripotta):

is this a multiple-choice question?

OpenStudy (omnomnom):

Um... Its not a multiple choice question

OpenStudy (aripotta):

dunno, sorry :/ i wish i could help

OpenStudy (omnomnom):

Aw... T_T The bad part is that i have to prove it too

OpenStudy (aripotta):

OHHHHHHHHH

OpenStudy (omnomnom):

OHHHHHH?

OpenStudy (aripotta):

ONE SEC.

OpenStudy (omnomnom):

Um... Okay.. O.o

OpenStudy (aripotta):

:/ nvm.

OpenStudy (aripotta):

i thought it was something, but it wasn't

OpenStudy (omnomnom):

URRGH

OpenStudy (aripotta):

sorry :l again. i'm failing today. and OMG. I FORGOT I WAS IN A TEST.

OpenStudy (omnomnom):

Oh lol xDD SMART!

OpenStudy (amistre64):

this is a system of 2 equations in 2 unknowns x + y = n 3x - 2y = n to solve this, i would use elimination; multiply the first by say; 2 2x + 2y = 2n 3x - 2y = n ------------- 5x + 0 = 3n solve for x; can do the same thing for y by multiplying the top by -3 and running thru the same steps

OpenStudy (omnomnom):

O.O wait what?

OpenStudy (omnomnom):

Isnt this simple algebra though... OMG.... O.O Its either that or x + y + 3x - 2y = 180 I can do either one because this was in an Geometry question with angles and stuff

hero (hero):

4x - y = 180

hero (hero):

How come you won't post the full question?

OpenStudy (omnomnom):

Urrgh idk... Fine i willl

OpenStudy (omnomnom):

But its a geometry questtion: RS intersects LM at P, measure Of angle RPL = x+y, measure of angle LPS = 3x +2y, Measure of angle MPS = 3x -2y . A. Solve for X and Y B. Find measure of angle RPL,LPS, AND MPS

hero (hero):

You need to draw the picture as well.

hero (hero):

Don't be lazy

OpenStudy (omnomnom):

Oh lol OKaaaaaaaaaay

hero (hero):

If you don't post the full question, you won't be able to get proper help

OpenStudy (omnomnom):



OpenStudy (aripotta):

omg. why didn't you post htis earlier.

OpenStudy (aripotta):

this* i could have helped you lol

OpenStudy (omnomnom):

Lol i didnt knooooooooooow

OpenStudy (aripotta):